# Search for the population of the low-energy continuum of tetraneutron system was performed for the H(He,Li) and H(He,He)HH+ reactions. Evidence for a hump

### 09 Jan 2024 v2

@@
H(He,Li reaction in the limit of highest H decay
energies. The population of the low-energy region in the n spectrum is
found to be perfectly correlated with the population of the lowest Li
@@ 1 Introduction
section of nb. Much more encouraging and intriguing was a very recent
@@
theoretical approaches Ref. [7, 8, 9]
always with the same result: for the two-body potentials considered fixed, a
radical modification of three-body potential is required, which is not tolerable
as leading to “devastating effect” on the nuclear structure in general. In the
@@
found for important low-energy effect of n final-state interaction (FSI).
These are: specific trajectories of the -matrix poles, enhanced “time
@@
to populate n continuum by the knockout of the remnant nuclides
@@
experimentally and interpreted theoretically as a few-body continuum response
combining effects of FSI and initial-state He structure (ISS) via the
@@
the strength function due to the peripheral character of both the halo wave
functions (WF) and electromagnetic transition operators
@@
the other reactions with He beam. As for the theoretical discussion we
revisit the n population from He in simple approach
@@ 2 Experiment
to be much more “treasurous” than intended, providing interesting results also
@@
The projectile trajectories were reconstructed by two pairs of multi-wire
@@
(PSD) [30]. The PSD information, augmented with ToF
@@ 2.1 H(He,Li)n results
compared with the reconstructed MM energy of the 4n
group. The fact that the majority of these events (108 events) are
located inside the kinematically allowed region proves the good channel
identification for the data. The empty-target measurement Fig. 2
(a) shows that the background conditions were very “clean” for the n
population in the H(He,Li) reaction: the “empty” events are very
few and located mainly in the unphysical part of the kinematical plane.
@@
events (black diamonds). The empty-target measurement events are shown by green
circles. (b) The n MM spectrum. Green dots show the
@@
the measured Li recoil energy and emission angle, taking the 108 events with
the data points located within the “kinematical triangle” . The obtained spectrum shows a group of 6 events around
MeV. The width of this group is well consistent with the
value of energy resolution MeV in this region of n MM spectrum.
The statistics of this low-energy group is marginal, but resolution is
sufficient and background conditions are shown above to be very good.
@@
terms of the hyperspherical harmonics (HH) model. The hypermomentum is grand
quantum number in the few-body configuration space. This leads to a modified
4-body phase volume behavior . It is notable that the observed
energy of this group events ( MeV) is consistent (within the
@@ 2.2 H(He,He)HH+ results
H(He,He)H data of [25]. Green dots show the
empty target events. Red dashed and blue dotted frames show the event selection
for Fig. 4 (b). (b) Tetraneutron MM spectrum summed up for
MeV. Large green circles show the simulated empty-target
spectrum. Thick gray line is MC simulated phase volume for H decays.
@@
H decay product of H, which allows to infer MM of the n
subsystem. These data in terms of the n emission are shown in Fig. 3 (a). The H events are mainly concentrated close to
diagonal , which means that the “internal” energy of
the tetraneutron tends to be large, while the triton gets only a small fraction
of the total H decay energy. This situation seems to be close to the phase
volume distribution of the 5-body H+ decay for given H decay
energy as
@@
value for 5-body phase volume is . It was discussed in the previous
section that at least value is requested for by the Pauli principle for n. This idea is confirmed by the
@@
distributions at given H energies are also quite “compressed” due to
experimental bias, see Fig. 13 of [25], so for
distribution Eq. (1) turns out to be completely justified in our
situation.
To observe the possible n correlation, we need to find the kinematical
region, where the main “phase volume” component of the decay distribution Eq. (1) is well suppressed. This means , and
@@
region. The background of about events/MeV (green dots in Fig. 3 (b) is estimated from the available empty target data. Thus
significant part of this low-energy n data may be connected with
the random background.
Fortunately, it appeared that the low-energy n events in Fig. 3 (b) are nicely correlated with the population of the lowest
MeV state in the He+H continuum of the
[32], see Fig. 4 (b).
This means that we are actually dealing here not with the HH+ decay, but rather with the H(He,Li reaction.
Moreover, the same correlation is perfectly true not only for events in the
MeV range, but also for the MeV range. In
this range most of the events should be either unphysical or connected with low
energy resolution, as the population of the MeV range is
predicted to be negligible in all theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy
@@
number of n events with MeV. The coincidence with
definite state in Li practically guarantee the physical nature of such
event, so we recognize that absolute majority (26 events out of 28) found in the
range MeV in Fig. 3 (b) as really
belonging to n MM spectrum.
@@
that “high-energy” parts of n spectrum are qualitatively different for
the ranges and MeV. This
may be indication that within these energy ranges there are actually two states
of Li: and MeV [32]. So far this
speculation should be taken with caution because of concentration of “empty
target” events in the high n MM energy part of these spectra, see Fig. 3 (a).
@@ 3 On n population in different reactions
channel for the reaction of Ref. [6] and reactions considered in
@@
reaction are shown by black circles in (a). The same for the
@@
red stars in (a) corresponds to ranges
and MeV, blue diamonds and red stars in (b) corresponds to
@@
where highly excited states of Li (above the He-H threshold)
are populated. We find in Fig. 4 (c) that the most expressed
low-energy n indication is obtained for the recoil,
which is the lowest He-H resonant state.
@@
intermediate energies, which correspond to + continuum states of the
@@
“soft” H projectile. So far, the possibility to extract this information
from our existing data is questionable.
@@ 4 Theoretical model
The easiest reaction model which can be considered for n production is
@@ 8 Conclusion
same type of phenomenon. The statistics of 6 events for the first reaction is
quite low, but events for the second reaction is comparable to
statistics connected with the low-energy structure observed in
[6].
@@
[19] and of this work. As compared to [19]
(i) the calculations of this work are performed with realistic NN and 3N
interactions, and (ii) the obtained realistic He source demonstrates the
complicated correlated behavior that is very important for the n
population. However, these two aspects are not sufficient to fully explain the
presence of a low-lying peak in the tetraneutron spectrum. To do this, it is
also necessary to consider extremely peripheral reaction mechanism, in which the
-core of He interacts with the target, but four neutrons remain
“spectators” at distances, far exceeding the typical sizes of neutron orbitals
in He. Using the appropriate profile function of the type
(10), simulating such a geometry, it is possible to explain the
presence of a low-lying peak in the tetraneutron spectrum. Since the appearance
of the low-energy peak is not related to the tetraneutron per se, but
also to the source used and the reaction mechanism, it should be expected that
it could be detected at somewhat different energies in different reactions. A
possible reason of disagreement with the calculations [23] may
be the incorrect n c.m. treatment in He WF constructed in
[23].

### 20 Jan 2024 v4

@@
H(He,Li reaction in the limit of the highest H
decay energies. The population of the low-energy region in the n spectrum
is found to be perfectly correlated with the population of the lowest Li
@@ 1 Introduction
section of nb. Much more encouraging and intriguing was very recent
@@
theoretical approaches [7, 8, 9] always
with the same result: for the two-body potentials considered fixed, a radical
modification of three-body potential is required, which is not tolerable as
leading to “devastating effect” on the nuclear structure in general. In the
@@
found for important low-energy effect of the n final-state interaction
(FSI). These are: specific trajectories of the -matrix poles, enhanced “time
@@
to populate n continuum by the knockout of the remnant nuclide
@@
experimentally and understood theoretically as a few-body continuum response
combining the effects of FSI and the initial-state He structure (ISS) via
the
@@
the strength function due to the peripheral character of both, the halo wave
functions (WF) and the electromagnetic transition operators
@@
the other reactions with the He beam. As for the theoretical discussion,
we revisit the n population from He in simple approach
@@ 2 Experiment
to be much “richer” than intended, providing interesting results also
@@
The projectile trajectories were reconstructed by the two pairs of multi-wire
@@
(PSD) [30]. The PSD information, supplemented with ToF
@@ 2.1 H(He,Li)n results
compared with the reconstructed MM energy of the group. The fact
that the majority of these events (108 events) are located inside the
kinematically allowed region proves a good channel identification for the data.
The empty-target measurement Fig. 2 (a) shows that the background
conditions were very “clean” for the n population in the
H(He,Li) reaction: the “empty” events are very few and located
mainly in the unphysical part of the kinematical plane.
@@
events (black diamonds). The empty-target measurement events are shown by large
green circles. (b) The n MM spectrum. Green doted curve show the
@@
the measured Li recoil energy and emission angle, taking the 108 events
located within the “kinematical triangle” . The obtained
spectrum shows a group of 6 events around MeV. The width of
this group is well consistent with the value of energy resolution MeV
FWHM in this region of n MM spectrum. The statistics of this low-energy
group is marginal, but the energy resolution is sufficient and the background
conditions are shown above to be very good.
@@
terms of the hyperspherical harmonics (HH) model. The hypermomentum is the
grand quantum number in the few-body configuration space. This leads to a
modified 4-body phase-volume behavior . It is notable that the
observed energy of this group events ( MeV) is consistent (within the
@@ 2.2 H(He,He)HH+ results
H(He,He)H data of [25]. Large green circles
show the empty target events. Red dashed and blue dotted frames show the event
selection for Fig. 4 (b). (b) Tetraneutron MM spectrum summed
up for MeV. Large green circles show the simulated
empty-target spectrum. Thick gray line is MC simulated phase volume for H
decays.
@@
H decay product of H, which allows one to infer the MM of the
n subsystem. These data in terms of the n emission are shown in
Fig. 3 (a). The H events are mainly concentrated close
to diagonal , which means that the “internal” energy
of the tetraneutron tends to be large, while the triton gets only a small
fraction of the total H decay energy. This situation seems to be close to
the phase-volume distribution of the 5-body H+ decay for the given
H decay energy as
@@
value for the 5-body phase volume is . It was discussed in the
previous section that at least value is requested for by the Pauli principle for n. This idea is confirmed by the
@@
distributions at given H energies are also quite
“compressed” due to the experimental bias, see Fig. 13 of
[25], so for the distribution of Eq. (1) turns out to be completely justified in our situation.
To observe a possible n correlation, we need to find the kinematical
region, where the main “phase-volume” component of the decay distribution Eq. (1) is well suppressed. This means , and
@@
region. The background of about events/MeV is estimated from the
available empty target data, see green circles in Fig. 3 (b).
Thus a significant part of this low-energy n data may be
connected with the random background.
The H data can be converted to presentation of
system reconstructed as He+H continuum, see Fig. 4
(a). Fortunately, it appeared that the low-energy n events in Fig. 3 (b) are nicely correlated with the population of the lowest
MeV [32] state in the
He+H continuum, see Fig. 4 (b). This means that we
are actually dealing here not with the HH+ decay, but
rather with the H(He,Li reaction. Moreover, the same
correlation is perfectly true not only for events in the MeV
range, but also for the MeV range. In this range most of the
events should be either unphysical or connected with insufficient energy
resolution, as the population of the MeV range is predicted
to be negligible in available theoretical scenarios, see Fig. 7 or Ref. [23]. The MC estimated energy
@@
number of events with MeV. The coincidence with the definite
state in Li practically guarantee the physical nature of such event, so we
recognize that the absolute majority of events (26 out of 28) found in the range
MeV in Fig. 3 (b) as really belonging to
the n MM spectrum.
@@
that “high-energy” parts of the n spectrum are qualitatively different
for the ranges and MeV.
This may be indication that within these energy ranges there are actually two
states of Li: and MeV [32].
@@ 3 On n population in different reactions
channel for the reaction of Ref. [6] and the reactions considered
in
@@
reaction is shown by black circles in (a). The same for the
@@
red stars in (a) correspond to ranges
and MeV, blue diamonds and red stars in (b) correspond to
@@
where highly excited states of Li (those located above the
He-H threshold) are populated. We find in Fig. 4
(c) that the most expressed low-energy n hump is obtained for the
recoil, which is the lowest He-H resonant
state.
@@
intermediate energies, which correspond to + continuum states of
@@
“soft” H target. So far, the possibility to extract this information from
our existing data is questionable.
@@ 4 Theoretical model
The simpliest reaction model which can be considered for n production is
@@ 8 Conclusion
same type of phenomenon. The statistics of 6 events for the first our reaction
is quite low, but events for the second one is comparable to statistics
connected with the low-energy structure observed in [6].
@@
[19] and the results of this work. As compared to
[19] the calculations of this work (i) are performed with
realistic NN and 3N interactions, and (ii) the obtained realistic He
source demonstrates the complicated correlated behavior that is very important
for the n population. However, these two aspects are not sufficient to
fully explain the presence of a low-lying peak in the tetraneutron spectrum. To
do this, it is also necessary to consider extremely peripheral reaction
mechanism, in which the -core of He interacts with the target, but
four neutrons remain “spectators” at distances, far exceeding the typical
sizes of neutron orbitals in He. Using the appropriate profile function of
the type Eq. (10), simulating such a geometry, it is possible to
explain the presence of a low-lying peak in the tetraneutron spectrum. Since the
appearance of the low-energy peak is not related to the tetraneutron per
se, but also to the ISS (the source used) and the reaction mechanism, it should
be expected that such a structure could be detected at somewhat different
energies in different reactions. A possible reason of disagreement with the
calculations [23] may be the incorrect n c.m. treatment
in the He WF constructed in [23].